1. solve 20= t/15
2. a/6= -108

Answer by mananth(12270)

20=t/15
multiply by 15
2*15=t/15 * 15
t=300
...
a/6=-108
multiply by 6
a/6 * 6 = -108*6
a=-648
...
m.ananth@hotmail.ca